OpenStudy (anonymous):

g(x)=6/2-3x find the domain of the function

OpenStudy (anonymous):

domain = real numbers except x = 2/3

OpenStudy (anonymous):

Your function really is f(x)=6/(2-3x)

OpenStudy (anonymous):

x\[\neq\]2/3 or is it x\[\ge\]2/3

jimthompson5910 (jim_thompson5910):

Domain is \[\large \left\{x|x\in\mathbb{R}, x\neq\frac{2}{3} \right\}\]

OpenStudy (anonymous):

The denominator can't equal zero; all other real number are ok. so set the denominator equal to 0 and exclude that number \[2-3x=0\] \[2=3x\] \[x=2/3\] The domain is all real numbers except 2/3

jimthompson5910 (jim_thompson5910):

basically, x can be any number but it cannot be 2/3

OpenStudy (anonymous):

In interval notation it would be \[(-\infty,2/3) or (2/3,\infty)\] Jim_t showed the set notation above

jimthompson5910 (jim_thompson5910):

The domain in interval notation is \[\large \left(-\infty, \frac{2}{3}\right)\cup\left(\frac{2}{3},\infty \right)\]